OpenStudy (anonymous):

Suppose a bottle of wine at a room temperature of 72 degree fahrenheit is place in a refrigerator at 40 degree fahrenheit to cool before a dinner party. After an hour the temperature of the wine is found to be 61.5 degree fahrenheit. Find the constant k, to two decimal places, and the time, to one decimal place it will take the wine to cool from 72 to 50 degree fahrenheit

OpenStudy (anonymous):

\[ T=T_{m}+(T_{0}-T_m)e^-kt\]

OpenStudy (anonymous):

that is the equation

OpenStudy (anonymous):

e is raise to -kt

OpenStudy (anonymous):

T=Tm+(To-Tm)=e^-kt 61.5=40 +(72-40)e^-k(1hour) (61.5-40)=(72-40)e^-k(1hour) (61.5-40)/(72-40)=e^-k(1hour) 21.5/32 =e^-k(1) taking ln on both sides gives ln(21.5/32)= -k -0.40=-k k=0.40 now 50=40+(72-40)e^-0.4t 50-40=(72-40)e^-0.4t 10=32e^-0.4t 10/32=e^-0.4t taking ln on both sides ln(10/32)=-0.40t t= -ln(10/32)/0.40 t=2.9 hours ...ans

OpenStudy (anonymous):

if you have any question let me know

OpenStudy (anonymous):

is everything ok? any question xiongvue?